### v3

# Every Poincaré gauge theory is conformal: a compelling case for dynamical vector torsion

Abstract
The Poincaré gauge theory (PGT) of gravity provides a viable formulation of general relativity (Einstein–Cartan theory), and a popular model-building framework for modified gravity with torsion. Notoriously, however, the PGT terms which propagate vector torsion lead to strongly-coupled ghosts: the modern view is that only scalar torsion can propagate. To fix this, we revisit the concept of embedding explicit mass scales in scale-invariant theories, showing how the Klein–Gordon theory naturally leads to a slowly-rolling inflaton. We then show that the unique scale-invariant embedding of PGT leads to two new terms, one of which is the Maxwell term for vector torsion. We provide the full spectrum of quantum particles in the resulting theory. Our result means that every PGT is conformal and – after a two-decade hiatus – vector torsion is back on the menu.
Introduction to torsion
General relativity (GR) offers a remarkably successful description of gravity as spacetime curvature [1]. [^1] The term ‘GR’ refers strictly to the textbook, metric-based theory with an Einstein–Hilbert action in Riemann spacetime. An open problem is the degree – if any – to which spacetime torsion also participates in the gravitational interaction. There are excellent theoretical reasons to consider torsion: it was shown by Utiyama [2], Kibble [3] and Sciama [4] that when fermions are coupled to gravity, the natural result is a local gauge theory not only of spacetime translations (i.e. the diffeomorphisms of GR), but also of spatial rotations and Lorentz boosts. In this Poincaré gauge theory (PGT) of gravity [5], the fundamental fields can be [^2] By contrast with GR, the term ‘PGT’ encompasses any action formulated with the 40 d.o.f of the tetrad and spin connection in Riemann–Cartan spacetime. In the absence of fermionic matter, the dynamics of PGT are also fully captured by the alternative post-Riemannian field parameterisation of 34 d.o.f in $\big{\{}\tensor{g}{{}_{\mu\nu}},\tensor{T}{{}^{\lambda}_{\mu\nu}}\big{\}}$, though the Poincaré symmetry is hidden. the metric tensor $\tensor{g}{{}_{\mu\nu}}$ and an a priori independent affine connection $\tensor{\Gamma}{{}^{\alpha}_{\mu}{}_{\nu}}$ which gives rise (as we will show later) to torsion and curvature as the field strength tensors of spacetime translations and rotations, respectively:
$\tensor{T}{{}^{\alpha}_{\mu}{}_{\nu}}$ $\equiv 2\tensor{\Gamma}{{}^{\alpha}_{[\mu}{}_{\nu]}},$ (1a)
$\tensor{R}{{}^{\rho}_{\sigma}{}_{\mu\nu}}$ $\equiv 2\left(\tensor{\partial}{{}_{[\mu|}}\tensor{\Gamma}{{}^{\rho}_{|\nu]}{}_{\sigma}}+\tensor{\Gamma}{{}^{\rho}_{[\mu|}{}_{\alpha}}\tensor{\Gamma}{{}^{\alpha}_{|\nu]}{}_{\sigma}}\right).$ (1b)
To connect this to textbook GR, if $\tensor{T}{{}^{\alpha}_{\mu}{}_{\nu}}\to 0$ in Eq. 1a then $\tensor{\Gamma}{{}^{\mu}_{\nu}{}_{\rho}}$ loses its independence from the metric [^3] We assume the metricity condition. and is fixed by the Christoffel formula $\tensor{\Gamma}{{}^{\mu}_{\nu}{}_{\rho}}\to\tensor{\mathring{\Gamma}}{{}^{\mu}_{\nu\rho}}\equiv\tensor{g}{{}^{\mu\lambda}}\big{(}\tensor{\partial}{{}_{(\nu}}\tensor{g}{{}_{\rho)\lambda}}-\frac{1}{2}\tensor{\partial}{{}_{\lambda}}\tensor{g}{{}_{\nu\rho}}\big{)}$. In this torsion-free limit, Eq. 1b defines the Riemannian curvature $\tensor{R}{{}^{\rho}_{\sigma}{}_{\mu\nu}}\to\tensor{\mathring{R}}{{}^{\rho}_{\sigma}{}_{\mu\nu}}$. Despite concrete theoretical provenance from PGT, there is uncertainty over what a signal for torsion might look like in the phenomena. This is in contrast to curvature, whose presence can be inferred e.g. through the geodesic motion of test particles. To understand this uncertainty, notice that we can always perform the field reparameterisation [^4] Note $\tensor{\Gamma}{{}^{\mu}_{\nu}{}_{\rho}}-\tensor{\mathring{\Gamma}}{{}^{\mu}_{\nu}{}_{\rho}}$ is sometimes called contortion, and is linear in torsion. $\tensor{\Gamma}{}\mapsto\tensor{\mathring{\Gamma}}{}+\tensor{T}{}$ (leading to $\tensor{R}{}\mapsto\tensor{\mathring{R}}{}+\tensor{\partial}{}\tensor{T}{}+\tensor{T}{}^{2}$ with indices suppressed). This post-Riemannian expansion from variables $\big{\{}\tensor{g}{{}_{\mu\nu}},\tensor{\Gamma}{{}^{\alpha}_{\mu}{}_{\nu}}\big{\}}$ to $\big{\{}\tensor{g}{{}_{\mu\nu}},\tensor{T}{{}^{\lambda}_{\mu\nu}}\big{\}}$ reveals all PGT models to be nothing more than metric-based gravity coupled – minimally or otherwise – to the field $\tensor{T}{{}^{\lambda}_{\mu\nu}}\equiv\tensor{T}{{}^{\lambda}_{[\mu\nu]}}$ as if it were an exotic kind of matter. Notice how there are infinitely many such models: if there are models for all experimental outcomes, then the theoretical framework ceases to be useful. The classic example of a useful PGT which establishes a predictive baseline is the minimal Einstein–Cartan theory (ECT) [6, 7, 8, 9, 10, 11, 12]. ECT is defined as the specific case of PGT which shares the Einstein–Hilbert action of GR
$S_{\text{ECT}}\equiv\int\mathrm{d}^{4}x\sqrt{-g}\Big{[}-\frac{{m_{\mathrm{P}}}{}^{2}}{2}\tensor{R}{}-{m_{\mathrm{P}}}{}^{2}\Lambda+\text{matter}\Big{]},$ (2)
where $\tensor{R}{}\equiv\tensor{R}{{}^{\mu}_{\mu}}\equiv\tensor{R}{{}^{\mu\nu}_{\mu\nu}}$ and ${m_{\mathrm{P}}}{}\equiv 1/\sqrt{\kappa}$ is the Planck mass, for $\kappa$ the Einstein constant. With $\Lambda\equiv 0$ and in the absence of matter, Eq. 2 implies the vacuum Einstein equations $\tensor{R}{{}_{\mu\nu}}=0$ and the auxiliary equation $\tensor{T}{{}^{\lambda}_{\mu\nu}}=0$. Inclusion of matter leads to a contact spin-torsion coupling; the torsion integrates out to leave effective four-Fermi interactions [3, 13, 14, 15, 16, 17, 18]. Such interactions have many pheonomenological applications [14, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 18, 55, 56, 57, 58], such as to fermionic dark matter production. Thus, non-propagating torsion might reasonably be constrained (albeit weakly) by dark matter abundances. Propagating torsion, on the other hand, can come from any number of next-to-minimal $\big{\{}\tensor{g}{{}_{\mu\nu}},\tensor{T}{{}^{\lambda}_{\mu\nu}}\big{\}}$ models: it could be sufficiently heavy to evade current bounds, or conveniently account for any new scalar or vector bosons that might be observed (note that $\tensor{T}{{}^{\lambda}_{\mu\nu}}$ contains various bosons up to spin-two). In other words, the phenomological richness of the full $\big{\{}\tensor{g}{{}_{\mu\nu}},\tensor{T}{{}^{\lambda}_{\mu\nu}}\big{\}}$ theory-space, which has infinitely many parameters, renders it non-predictive.
No-go theorem for vector torsion
Happily, the literature has a convincing way to deal with this problem. Just as PGT motivates torsion, so the usually assumed structure of its action also endows torsionful model-building with predictivity. The logic is that PGT should be restricted (see e.g. [59, 60, 61, 62, 63]) to a Yang–Mills-type action
$S_{\text{PGT}}\equiv\int\mathrm{d}^{4}x\sqrt{-g}\Big{[}-\frac{{m_{\mathrm{P}}}^{2}}{2}\tensor{R}{}-{m_{\mathrm{P}}}{}^{2}\Lambda+\tensor{\alpha}{{}_{1}}\tensor{R}{}^{2}+\tensor{R}{{}_{\mu\nu}}\big{(}\tensor{\alpha}{{}_{2}}\tensor{R}{{}^{\mu\nu}}$
$+\tensor{\alpha}{{}_{3}}\tensor{R}{{}^{\nu\mu}}\big{)}+\tensor{R}{{}_{\mu\nu\sigma\lambda}}\big{(}\tensor{\alpha}{{}_{4}}\tensor{R}{{}^{\mu\nu\sigma\lambda}}+\tensor{\alpha}{{}_{5}}\tensor{R}{{}^{\mu\sigma\nu\lambda}}+\tensor{\alpha}{{}_{6}}\tensor{R}{{}^{\sigma\lambda\mu\nu}}\big{)}$
$+\tensor{T}{{}_{\mu\nu\sigma}}\big{(}\tensor{\beta}{{}_{1}}\tensor{T}{{}^{\mu\nu\sigma}}+\tensor{\beta}{{}_{2}}\tensor{T}{{}^{\nu\mu\sigma}}\big{)}+\tensor{\beta}{{}_{3}}\tensor{T}{{}_{\mu}}\tensor{T}{{}^{\mu}}+\text{matter}\Big{]},$ (3)
where $\tensor{T}{{}_{\mu}}\equiv\tensor{T}{{}^{\nu}_{\nu\mu}}$, and No-go theorem for vector torsion extends Eq. 2 by only admitting quadratic field strength invariants. [^5] In this letter we omit the parity-odd invariants only out of simplicity; there are no convincing theoretical grounds for excluding them [64, 65]. Evidently, No-go theorem for vector torsion occupies just a small corner of the $\big{\{}\tensor{g}{{}_{\mu\nu}},\tensor{T}{{}^{\lambda}_{\mu\nu}}\big{\}}$ theory-space: the Yang–Mills structure is similar to the gauge boson sector of the standard model (SM), and it imposes surprising limits on the phenomenology. Firstly, the parameters in No-go theorem for vector torsion must be carefully tuned to eliminate ghosts and tachyons from the linearised spectrum [66, 67, 68, 69, 61, 62, 5, 63, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 64, 81, 82, 83, 84, 85, 65, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107]. Even then, inconsistencies usually reappear at non-linear order due to strongly coupled particles [108, 109, 59, 110, 111, 111, 112, 113, 114, 115, 116, 62, 117, 118, 119, 120, 121, 122, 122, 123, 124, 125, 126, 127, 128, 129, 130, 103, 129, 55, 131] (for general strong coupling see [132, 133, 134, 119, 120, 121, 135, 136, 125]) and the breaking of so-called ‘accidental’ symmetries which only survive linearly [137, 138, 139, 140, 141, 61, 62, 142, 143, 144, 145, 146, 130] (see also [60, 67, 68, 62, 5, 79, 147, 148, 65, 29, 149, 98]). Weakly coupled, consistent PGTs are very rare: the current consensus is that only spin-zero scalar torsion is allowed to propagate, and then only with very specific non-linear interactions [140, 141, 61, 62] (see e.g. [70, 71, 74, 76, 77, 78, 82, 86, 89, 91, 97, 101] for applications, [63, 75, 80, 81, 95] for reviews, and [88, 150, 151, 96, 99, 152, 100, 153, 154, 155] for similar analyses of spacetime non-metricity). Spin-one vector torsion, on the other hand, is ruled out [61, 62]. Purely vector torsion can only be linearly propagated by the $\tensor{R}{{}_{[\mu\nu]}}\tensor{R}{{}^{[\mu\nu]}}$ term, reached by setting $\tensor{\alpha}{{}_{2}}+\tensor{\alpha}{{}_{3}}=0$ in No-go theorem for vector torsion. However, that term carries in strongly-coupled ghosts, so the resulting theory is inconsistent [61, 62]. In a recent attempt to evade this no-go theorem, we introduced multiplier fields which pacify this strong coupling [55]. A fair criticism of our approach, however, was that it effectively reduced $\tensor{R}{{}_{[\mu\nu]}}\tensor{R}{{}^{[\mu\nu]}}$ to the post-Riemannian Maxwell term $\tensor{\partial}{{}_{[\mu}}\tensor{T}{{}_{\nu]}}\tensor{\partial}{{}^{[\mu}}\tensor{T}{{}^{\nu]}}$ which – though non-linearly consistent – is absent from No-go theorem for vector torsion. This Maxwell term can still be added by hand, as was done for example in [104]. The problem with adding terms manually is that, since almost any desired dynamics can then be fabricated, the resulting models are less predictive and compelling. To summarise: every PGT in No-go theorem for vector torsion has a post-Riemannian expansion, but not every covariant model of $\big{\{}\tensor{g}{{}_{\mu\nu}},\tensor{T}{{}^{\lambda}_{\mu\nu}}\big{\}}$ is the post-Riemannian expansion of a PGT in No-go theorem for vector torsion.
In this letter
We show that the $\tensor{\partial}{{}_{[\mu}}\tensor{T}{{}_{\nu]}}\tensor{\partial}{{}^{[\mu}}\tensor{T}{{}^{\nu]}}$ term is a Yang–Mills-type term in disguise, without which No-go theorem for vector torsion is actually incomplete. Our reasoning is that PGT is not the most fundamental gauge theory of gravity. Recently we developed an extended [^6] Given the considerations to be presented in this letter, perhaps it would be more appropriate to replace ‘extended’ with ‘economical’. Nonetheless, we retain the original nomenclature for consistency with the existing literature. version of Weyl gauge theory (eWGT), which gauges the full conformal group [156, 157, 158, 159, 160]. We will show that the covariant derivative of eWGT reduces precisely to that of PGT when it is expressed in terms of scale-invariant variables (equivalent to fixing the scale gauge, and so breaking the conformal symmetry to Poincaré symmetry). The Yang–Mills-type action of eWGT reduces to No-go theorem for vector torsion plus new terms (to be shown in Equivalence of PGT and eWGT). This provides a concrete motivation for vector torsion, which was hitherto lacking. Before obtaining these results, we introduce scale-invariant embeddings with a simple toy model, which is nonetheless interesting.
Figure 1: Massive scalar perturbations in Eq. 4 have a unique, non-linear, scale-invariant embedding given in Scale-invariant embedding. This embedding can lead to only one qualitative alteration in Eqs. 8a and 8b: the quadratic potential (dashed, with a unit-circle illustrating the mass scale $m$) develops a non-linear plateau (cyan) whose depth and height depend on model parameters $\nu$ and $\sigma$. The plateau is consistent with current CMB constraints on a slowly-rolling inflaton (arrow).
Scale-invariant embedding
For our toy model, we work just with the matter sector. Let us suppose we know that some scalar perturbations $\delta\varphi$ have a Klein–Gordon mass $m$
$S=\int\mathrm{d}^{4}x\sqrt{-g}\left[{\textstyle\frac{1}{2}}\tensor{\partial}{{}_{\mu}}\delta\varphi\ \tensor{\partial}{{}^{\mu}}\delta\varphi-{\textstyle\frac{1}{2}}m^{2}\delta\varphi^{2}+\text{gravity}\right],$ (4)
where we allow the theory to be minimally coupled to a metric-based gravity theory (e.g. GR). This situation already arises in cosmology: we typically infer that the canonical inflaton potential is approximated by a mass term $U(\varphi)=m^{2}\varphi^{2}/2+\dots$ in order to facilitate reheating into the SM plasma after the end of inflation [^7] Strictly, this is only true for ‘textbook’ slow-roll inflation [161, 162, 163]. We particularly note that Higgs inflation does not require a reheating mass [164]. [165]. This inference does not, however, place constraints on $U(\varphi)$ beyond the $\delta\varphi$ linear regime. For a host of reasons the most favoured non-linear completion envisages $\varphi$ rolling off a plateau in $U(\varphi)$ sufficiently slowly to drive $N\sim 55$ e-folds of inflation [166, 167, 168]. Quite how this plateau should be motivated through fundamental theory is not yet settled. What we want to show here is that, starting purely from the linear model in Eq. 4, there is a way to build up the non-linear plateau by asking that $m$ be a dynamically acquired mass scale in a locally scale-invariant ‘embedding’ theory (see Fig. 1). There is precedent here, since Eq. 4 also describes pions [169, 170, 171, 172]. Pion masses, however, emerge long after reheating, through chiral and electroweak symmetry breaking [173, 174]. Moreover, the absence of explicit mass scales only makes the SM globally scale-invariant. Under local rescalings, the metric deforms as $\tensor{g}{{}_{\mu\nu}}\mapsto e^{2\rho}\tensor{g}{{}_{\mu\nu}}$ and the scalar as $\varphi\mapsto e^{-\rho}\varphi$ for local $\rho\equiv\rho(x)$. Considering the first term in Eq. 4, we must therefore introduce a Weyl vector $\tensor{B}{{}_{\mu}}\mapsto\tensor{B}{{}_{\mu}}-\tensor{\partial}{{}_{\mu}}\rho$ to define a new covariant derivative on scalars $\tensor{\mathscr{D}}{{}_{\mu}}\varphi\equiv\big{(}\tensor{\partial}{{}_{\mu}}-\tensor{B}{{}_{\mu}}\big{)}\varphi$. The kinematic d.o.f are thus increased from $\big{\{}\tensor{g}{{}_{\mu\nu}},\varphi\big{\}}$ to $\big{\{}\tensor{g}{{}_{\mu\nu}},\varphi,\tensor{B}{{}_{\mu}}\big{\}}$. Next, considering the second term in Eq. 4, we must introduce a compensator scalar $\phi\mapsto e^{-\rho}\phi$ to make up the mass dimension of $m$, which now becomes a dimensionless parameter. In the effective field theory approach, the non-linear completion of Eq. 4 includes all scale-invariant operators of mass-dimension four in the new variables $\big{\{}\tensor{g}{{}_{\mu\nu}},\varphi,\tensor{B}{{}_{\mu}},\phi\big{\}}$, namely
$S=\int\mathrm{d}^{4}x\sqrt{-g}\Big{[}$ $\frac{1}{2}\tensor{\mathscr{D}}{{}_{\mu}}\varphi\tensor{\mathscr{D}}{{}^{\mu}}\varphi-\frac{\sigma}{2}\tensor{\mathscr{D}}{{}_{\mu}}\varphi\tensor{\mathscr{D}}{{}^{\mu}}\phi+\frac{\nu}{2}\tensor{\mathscr{D}}{{}_{\mu}}\phi\tensor{\mathscr{D}}{{}^{\mu}}\phi$
$-\frac{\mu^{2}}{2}\phi^{2}\varphi^{2}-\frac{\xi}{16}\tensor{H}{{}_{\mu\nu}}\tensor{H}{{}^{\mu\nu}}+\text{gravity}\Big{]},$ (5)
where $\mu$$\sigma$$\nu$ and $\xi$ are dimensionless model parameters, and $\tensor{H}{{}_{\mu\nu}}\equiv 2\tensor{\partial}{{}_{[\mu}}\tensor{B}{{}_{\nu]}}$ is the natural Weyl field-strength tensor, i.e. $\tensor{\mathscr{D}}{{}_{[\mu}}\tensor{\mathscr{D}}{{}_{\nu]}}\varphi\equiv-\frac{1}{2}\tensor{H}{{}_{\mu\nu}}\varphi$. For brevity we omit the $\varphi^{4}$ and $\phi^{4}$ self-interaction terms, though we still expect inflation to be a viable outcome with these terms present. [^8] We will find that the $\varphi^{4}$ operator deforms the potential without (visually) compromising the desired slow-roll plateau, though a detailed comparison with the CMB is lacking; the $\phi^{4}$ operator provides a cosmological constant, which can be neglected when considering the early-Universe physics.
Scale-invariant variables
One can compare $\tensor{B}{{}_{\mu}}$ to the electromagnetic four-vector potential. But unlike in scalar electrodynamics, where the matter fields carry a non-physical phase as well as a physical magnitude, the fields $\varphi$ and $\phi$ each have only one real d.o.f which merges entirely with the choice of scale gauge. This means that the field equations of Scale-invariant embedding cannot simultaneously propagate $\varphi$ and $\phi$ from the initial data: [^9] The scale-invariance of this model is local, i.e. it constitutes a gauge symmetry. To propagate the fields in the presence of a gauge symmetry, not only must the initial data be specified at some time-slice, but also the choice of gauge (in our case, the value of $\phi$) at all future times must be agreed upon [175]. The same scenario arises in classical electromagnetism and in GR: it does not detract from the well-posedness of the model. we must first declare a global solution for one of these fields. Remembering that we want to recover Eq. 4, we may take the opportunity to eliminate $\phi$ by assuming constant $\phi=\phi_{0}$. This is the Einstein choice of scale gauge. Rather than imposing the Einstein gauge, it is instead possible to eliminate $\phi$ completely from Scale-invariant embedding through the field reparameterisations $\tensor{g}{{}_{\mu\nu}}\mapsto(\phi_{0}/\phi)^{2}\tensor{\hat{g}}{{}_{\mu\nu}}$, $\varphi\mapsto\phi\hat{\varphi}/\phi_{0}$ and $\tensor{B}{{}_{\mu}}\mapsto\tensor{\hat{B}}{{}_{\mu}}+\tensor{\partial}{{}_{\mu}}\ln\phi$, resulting in
$S=\int\mathrm{d}^{4}x\sqrt{-\hat{g}}\Big{[}$ $\frac{1}{2}\left(\tensor{\partial}{{}_{\mu}}\hat{\varphi}-\left(\hat{\varphi}-\sigma\phi_{0}\right)\tensor{\hat{B}}{{}_{\mu}}\right)\left(\tensor{\partial}{{}^{\mu}}\hat{\varphi}-\hat{\varphi}\tensor{\hat{B}}{{}^{\mu}}\right)$
$-\frac{\mu^{2}\phi_{0}^{2}}{2}\hat{\varphi}^{2}-\frac{\xi}{16}\tensor{\hat{H}}{{}_{\mu\nu}}\tensor{\hat{H}}{{}^{\mu\nu}}+\frac{\nu\phi_{0}^{2}}{2}\tensor{\hat{B}}{{}_{\mu}}\tensor{\hat{B}}{{}^{\mu}}+\text{gravity}\Big{]},$ (6)
where we note $\tensor{\hat{H}}{{}_{\mu\nu}}\equiv 2\tensor{\partial}{{}_{[\mu}}\tensor{\hat{B}}{{}_{\nu]}}\equiv 2\tensor{\partial}{{}_{[\mu}}\tensor{B}{{}_{\nu]}}\equiv\tensor{H}{{}_{\mu\nu}}$. By construction, the new fields $\big{\{}\tensor{\hat{g}}{{}_{\mu\nu}},\hat{\varphi},\tensor{\hat{B}}{{}_{\mu}}\big{\}}$ are scale-invariant variables. The gauge symmetry in Scale-invariant variables seems to be destroyed, but we now understand that the rescaling transformations are still happening internally within the new fields. Scale-invariant variables and the Einstein gauge choice both result in Scale-invariant variables, a key feature of which is that the dimensionless couplings are now accompanied by powers of $\phi_{0}$. We will return to the natural values of these couplings later.
Bootstrapping inflation
After dropping the hats, shifting the vector by the scalar gradient to remove the cross-term and rescaling both fields [^10] Note during the rescaling $\tensor{B}{{}_{\mu}}\mapsto\tensor{B}{{}_{\mu}}/\sqrt{\xi}$ that unitarity requires $\xi\geq 0$., and for perturbations $\delta\varphi$ and $\delta\tensor{B}{{}_{\mu}}$ around the vacuum $\varphi=\tensor{B}{{}_{\mu}}=0$, we find that Scale-invariant variables reduces to
$S$ $=\int\mathrm{d}^{4}x\sqrt{-g}\Big{[}\frac{1}{2}\tensor{\partial}{{}_{\mu}}\delta\varphi\tensor{\partial}{{}^{\mu}}\delta\varphi-\frac{m^{2}}{2}\delta\varphi^{2}$
$-\frac{1}{4}\tensor{\partial}{{}_{[\mu}}\delta\tensor{B}{{}_{\nu]}}\tensor{\partial}{{}^{[\mu}}\delta\tensor{B}{{}^{\nu]}}+\frac{M^{2}}{2}\delta\tensor{B}{{}_{\mu}}\delta\tensor{B}{{}^{\mu}}+\text{gravity}\Big{]},$ (7)
[…]
$S=\int\mathrm{d}^{4}x\sqrt{-g}\left[{\textstyle\frac{1}{2}}\tensor{\partial}{{}_{\mu}}\varphi\tensor{\partial}{{}^{\mu}}\varphi-U\left(\varphi\right)+\text{gravity}\right],$ (8a)
$U\left(\varphi\right)\equiv\frac{\mu^{2}\phi_{0}^{4}}{2}\left[\frac{\sigma}{2}+\sqrt{\nu-\frac{\sigma^{2}}{4}}\text{sinh}\left(\frac{(\varphi-c)/\phi_{0}}{\sqrt{\nu-\frac{\sigma^{2}}{4}}}\right)\right]^{\ 2},$ (8b)
[…]
Recap and extension
Collecting our results so far:
  1. 1.
    Gravitational coupling to fermions naturally leads to PGT as a leading formulation of gravity. PGT does not have local scale invariance.
  2. 2.
    The $\tensor{\partial}{{}_{[\mu}}\tensor{T}{{}_{\nu]}}\tensor{\partial}{{}^{[\mu}}\tensor{T}{{}^{\nu]}}$ term for vector torsion $\tensor{T}{{}_{\mu}}$ is notoriously absent from the PGT action in No-go theorem for vector torsion.
  3. 3.
    Theories with mass scales have locally-scale-invariant embeddings which can lead to excellent phenomenology (inflation) besides a new massive vector $\tensor{B}{{}_{\mu}}$.
We will now connect these three observations by showing that eWGT is the unique scale-invariant embedding of PGT. This will identify $\tensor{T}{{}_{\mu}}/3$ with the vector $\tensor{B}{{}_{\mu}}$ when expressed in scale-invariant variables, and thereby reveal $\tensor{\partial}{{}_{[\mu}}\tensor{T}{{}_{\nu]}}\tensor{\partial}{{}^{[\mu}}\tensor{T}{{}^{\nu]}}$ to be a Yang–Mills-type term.
Poincaré gauge theory (PGT)
Gauge theories of gravity can be formulated on completely flat Minkowski spacetime, just like the SM [156]. In this formulation, we use holonomic Greek indices and Lorentz Roman indices. Translational gauge fields $\tensor{b}{{}^{i}_{\mu}}$ and inverses $\tensor{h}{{}_{i}^{\mu}}$ obey $\tensor{b}{{}^{i}_{\mu}}\tensor{h}{{}_{i}^{\nu}}\equiv\tensor{*}{\delta}{{}_{\mu}^{\nu}}$ and $\tensor{b}{{}^{i}_{\mu}}\tensor{h}{{}_{j}^{\mu}}\equiv\tensor{*}{\delta}{{}_{j}^{i}}$, and give rise to the curved-space metric $\tensor{g}{{}_{\mu\nu}}$ through $\tensor{b}{{}^{i}_{\mu}}\tensor{b}{{}^{j}_{\nu}}\tensor{\eta}{{}_{ij}}\equiv\tensor{g}{{}_{\mu\nu}}$. We assume contraction with these fields eliminates all coordinate indices of general (i.e. not necessarily scalar) matter fields $\varphi$, which have only suppressed Lorentz (or spinor) indices. A general $\varphi$ with Weyl weight $w$ and belonging to the $\mathrm{SL}(2,\mathbb{C})$ representation with Lorentz generators $\tensor{\Sigma}{{}_{ij}}\equiv\tensor{\Sigma}{{}_{[ij]}}$ transforms as $\varphi\mapsto e^{w\rho+\tensor{\omega}{{}^{ij}}\tensor{\Sigma}{{}_{ij}}/2}\varphi$, where $\tensor{\omega}{{}^{ij}}\equiv\tensor{\omega}{{}^{[ij]}}(x)$ are finite angles. The rotational gauge field is $\tensor{A}{{}^{ij}_{\mu}}\equiv\tensor{A}{{}^{[ij]}_{\mu}}$. The transformations $\tensor{b}{{}^{i}_{\mu}}\mapsto e^{\rho}\tensor{\Lambda}{{}^{i}_{j}}\tensor{b}{{}^{j}_{\mu}}$ and $\tensor{A}{{}^{ij}_{\mu}}\mapsto\tensor{\Lambda}{{}^{i}_{k}}\tensor{\Lambda}{{}^{j}_{l}}\tensor{A}{{}^{kl}_{\mu}}-\tensor{\Lambda}{{}^{jk}}\tensor{\partial}{{}_{\mu}}\tensor{\Lambda}{{}^{i}_{k}}$ (where $\tensor{\Lambda}{}\equiv e^{\tensor{\omega}{}}$ is the Lorentz matrix) ensure that the derivative $\tensor{\mathscr{D}}{{}_{i}}\equiv\tensor{h}{{}_{i}^{\mu}}\tensor{\mathscr{D}}{{}_{\mu}}$ with
$\tensor{\mathscr{D}}{{}_{\mu}}\varphi\equiv\left(\tensor{\partial}{{}_{\mu}}+{\textstyle\frac{1}{2}}\tensor{A}{{}^{ij}_{\mu}}\tensor{\Sigma}{{}_{ij}}\right)\varphi,$ (9)
is Lorentz-covariant. The field strengths from the commutator $\tensor{b}{{}^{[i|}_{\mu}}\tensor{b}{{}^{|j]}_{\nu}}\tensor{\mathscr{D}}{{}_{[i}}\tensor{\mathscr{D}}{{}_{j]}}\varphi\equiv\frac{1}{4}\tensor{b}{{}^{i}_{\sigma}}\tensor{b}{{}^{j}_{\lambda}}\tensor{R}{{}^{\sigma\lambda}_{\mu\nu}}\tensor{\Sigma}{{}_{ij}}\varphi-\frac{1}{2}\tensor{T}{{}^{\lambda}_{\mu\nu}}\tensor{\mathscr{D}}{{}_{\lambda}}\varphi$ are
$\tensor{T}{{}^{\lambda}_{\mu\nu}}$ $\equiv 2\tensor{h}{{}_{i}^{\lambda}}(\tensor{\partial}{{}_{[\mu|}}\tensor{b}{{}^{i}_{|\nu]}}+\tensor{A}{{}^{i}_{j}{}_{[\mu|}}\tensor{b}{{}^{j}_{|\nu]}}),$ (10a)
$\tensor{R}{{}^{\rho\sigma}_{\mu\nu}}$ $\equiv 2\tensor{h}{{}_{i}^{\rho}}\tensor{h}{{}_{j}^{\sigma}}(\tensor{\partial}{{}_{[\mu|}}\tensor{A}{{}^{ij}_{|\nu]}}+\tensor{A}{{}^{i}_{k}{}_{[\mu|}}\tensor{A}{{}^{kj}_{|\nu]}}),$ (10b)
where Eqs. 10a and 10b are equivalent to Eqs. 1a and 1b; the reparameterisation from gauge-theoretic $\big{\{}\tensor{b}{{}^{i}_{\mu}},\tensor{A}{{}^{ij}_{\mu}}\big{\}}$ to geometric $\big{\{}\tensor{g}{{}_{\mu\nu}},\tensor{\Gamma}{{}^{\alpha}_{\mu}{}_{\nu}}\big{\}}$ is described in [5, 156].
Weyl gauge theory (WGT)
From our toy model, the Weyl-covariant extension of Eq. 9 is
$\tensor{\mathscr{D}}{{}_{\mu}}\varphi\equiv\left(\tensor{\partial}{{}_{\mu}}+{\textstyle\frac{1}{2}}\tensor{A}{{}^{ij}_{\mu}}\tensor{\Sigma}{{}_{ij}}+w\tensor{B}{{}_{\mu}}\right)\varphi,$ (11)
and this construction defines the well-known Weyl gauge theory (WGT) of gravity [185, 186]. Whilst $\tensor{R}{{}^{\rho\sigma}_{\mu\nu}}$ in Eq. 10b is already Weyl-covariant, we have under local dilations
$\tensor{T}{{}^{\lambda}_{\mu\nu}}\mapsto\tensor{T}{{}^{\lambda}_{\mu\nu}}-2\tensor{*}{\delta}{{}^{\lambda}_{[\mu}}\tensor{\partial}{{}_{\nu]}}\rho\implies\tensor{T}{{}_{\mu}}\mapsto\tensor{T}{{}_{\mu}}-3\tensor{\partial}{{}_{\mu}}\rho,$ (12)
so from Eq. 12 only the combination
$\tensor{\tilde{T}}{{}^{\lambda}_{\mu\nu}}\equiv\tensor{T}{{}^{\lambda}_{\mu\nu}}-2\tensor{*}{\delta}{{}^{\lambda}_{[\mu}}\tensor{B}{{}_{\nu]}},$ (13)
is covariant, indeed $\tensor{b}{{}^{[i|}_{\mu}}\tensor{b}{{}^{|j]}_{\nu}}\tensor{\mathscr{D}}{{}_{[i}}\tensor{\mathscr{D}}{{}_{j]}}\varphi\equiv\frac{1}{4}\tensor{b}{{}^{i}_{\sigma}}\tensor{b}{{}^{j}_{\lambda}}\tensor{R}{{}^{\sigma\lambda}_{\mu\nu}}\tensor{\Sigma}{{}_{ij}}\varphi-\frac{1}{2}\tensor{\tilde{T}}{{}^{\lambda}_{\mu\nu}}\tensor{\mathscr{D}}{{}_{\lambda}}\varphi+\frac{1}{2}w\tensor{H}{{}_{\mu\nu}}\varphi$. Following Scale-invariant embedding, the parity-even Yang–Mills-type embedding of No-go theorem for vector torsion must then be strictly
$S_{\text{WGT}}\equiv\int\mathrm{d}^{4}x\sqrt{-g}\Big{[}-\frac{\zeta\phi^{2}}{2}\tensor{R}{}-\lambda\phi^{4}+\tensor{\alpha}{{}_{1}}\tensor{R}{}^{2}+\tensor{R}{{}_{\mu\nu}}\big{(}\tensor{\alpha}{{}_{2}}\tensor{R}{{}^{\mu\nu}}$
$+\tensor{\alpha}{{}_{3}}\tensor{R}{{}^{\nu\mu}}\big{)}+\tensor{R}{{}_{\mu\nu\sigma\lambda}}\big{(}\tensor{\alpha}{{}_{4}}\tensor{R}{{}^{\mu\nu\sigma\lambda}}+\tensor{\alpha}{{}_{5}}\tensor{R}{{}^{\mu\sigma\nu\lambda}}+\tensor{\alpha}{{}_{6}}\tensor{R}{{}^{\sigma\lambda\mu\nu}}\big{)}$
$+\phi^{2}\tensor{\tilde{T}}{{}_{\mu\nu\sigma}}\big{(}\tensor{\gamma}{{}_{1}}\tensor{\tilde{T}}{{}^{\mu\nu\sigma}}+\tensor{\gamma}{{}_{2}}\tensor{\tilde{T}}{{}^{\nu\mu\sigma}}\big{)}+\phi^{2}\tensor{\gamma}{{}_{3}}\tensor{\tilde{T}}{{}_{\mu}}\tensor{\tilde{T}}{{}^{\mu}}+\frac{\nu}{2}\tensor{\mathscr{D}}{{}_{\mu}}\phi\tensor{\mathscr{D}}{{}^{\mu}}\phi$
$-\frac{\xi}{16}\tensor{H}{{}_{\mu\nu}}\tensor{H}{{}^{\mu\nu}}+\frac{\chi}{2}\tensor{R}{{}_{\mu\nu}}\tensor{H}{{}^{\mu\nu}}+\text{matter}\Big{]},$ (14)
with dimensionless $\zeta$$\lambda$$\tensor{\gamma}{{}_{1}}$$\tensor{\gamma}{{}_{2}}$$\tensor{\gamma}{{}_{3}}$$\nu$$\xi$$\chi$ and $\alpha_{1}$ through $\alpha_{6}$. Crucially, everything is quadratic in field strengths.
Extended Weyl gauge theory (eWGT)
In moving from PGT to WGT, we have enlarged the kinematic space from $\big{\{}\tensor{b}{{}^{i}_{\mu}},\tensor{A}{{}^{ij}_{\mu}}\big{\}}$ to $\big{\{}\tensor{b}{{}^{i}_{\mu}},\tensor{A}{{}^{ij}_{\mu}},\tensor{B}{{}_{\mu}},\phi\big{\}}$, just like in our toy model. Unlike for our toy model, however, the enlargement proves unnecessary: it is more economical to recycle $\tensor{B}{{}_{\mu}}$ from within the rotational gauge field. Let the traceless part of the latter field be $\tensor{\tilde{A}}{{}^{ij}_{\mu}}\equiv\tensor{\tilde{A}}{{}^{[ij]}_{\mu}}$ with only 20 d.o.f, such that $\tensor{h}{{}_{j}^{\nu}}\tensor{\tilde{A}}{{}^{ij}_{\nu}}\equiv 0$. Then, using $\big{\{}\tensor{b}{{}^{i}_{\mu}},\tensor{\tilde{A}}{{}^{ij}_{\mu}},\tensor{B}{{}_{\mu}},\phi\big{\}}$ there is one (and only one) way to emulate the combined Poincaré and Weyl covariance of Eq. 11, namely
$\tensor{\mathscr{D}}{{}_{\mu}}\varphi\equiv\Big{(}\tensor{\partial}{{}_{\mu}}+{\textstyle\frac{1}{2}}\Big{[}$ $\tensor{\tilde{A}}{{}^{ij}_{\mu}}-{\textstyle\frac{2}{3}}\tensor{b}{{}^{[i|}_{\mu}}\tensor{h}{{}^{|j]}{}^{\nu}}\big{(}3\tensor{B}{{}_{\nu}}$
$\ \ \ -2\tensor{h}{{}_{k}^{\lambda}}\tensor{\partial}{{}_{[\lambda|}}\tensor{b}{{}^{k}_{|\nu]}}\big{)}\Big{]}\Sigma{{}_{ij}}+w\tensor{B}{{}_{\mu}}\Big{)}\varphi.$ (15)
Although unique, [^14] The extra terms in the square brackets in Eq. 15 have been constructed so as to emulate the inhomogeneous Lorentz transformation of the (missing) trace vector $\tensor{h}{{}_{j}^{\nu}}\tensor{A}{{}^{ij}_{\nu}}$, without simultaneously generating an (unwanted) inhomogeneous Weyl transformation. There is only one such construction. One might ask whether a similar trick can be used to replace the axial vector $\tensor{\epsilon}{{}^{k}_{ijl}}\tensor{h}{{}_{k}^{\nu}}\tensor{A}{{}^{ij}_{\nu}}$. The answer is of course affirmative, but the construction in this case does not need (and must not contain) $\tensor{B}{{}_{\mu}}$ because the axial vector and purely tensor parts of the Ricci rotation coefficients are already Weyl-covariant: only the trace vector part has an inhomogeneous Weyl transformation, which requires correcting via $\tensor{B}{{}_{\mu}}$ in the manner proposed. By carefully propagating this observation through to our main result in Equivalence of PGT and eWGT, we conclude that local scale invariance always leads to dynamical vector torsion, and not axial vector torsion.we will soon show that Eq. 15 is just one notational way to write the eWGT covariant derivative from [156]. Because eWGT shares the symmetries of WGT, it must also share the Yang–Mills-type action in Weyl gauge theory (WGT). The difference is that we must replace $\tensor{A}{{}^{ij}_{\mu}}$, as when going from Eq. 11 to Eq. 15. After this change, we notice from Eq. 13 that $\tensor{\tilde{T}}{{}_{\mu}}\equiv 0$ identically. Thus, $\tensor{\gamma}{{}_{3}}$ in Weyl gauge theory (WGT) drops out of eWGT.
Equivalence of PGT and eWGT
With the field reparameterisation $\tensor{B}{{}_{\mu}}\mapsto\tensor{V}{{}_{\mu}}+\frac{2}{3}\tensor{h}{{}_{k}^{\lambda}}\tensor{\partial}{{}_{[\lambda|}}\tensor{b}{{}^{k}_{|\mu]}}$ we find Eq. 15 is
$\tensor{\mathscr{D}}{{}_{\mu}}\varphi\equiv\Big{(}\tensor{\partial}{{}_{\mu}}+{\textstyle\frac{1}{2}}\Big{[}$ $\tensor{\tilde{A}}{{}^{ij}_{\mu}}-2\tensor{b}{{}^{[i|}_{\mu}}\tensor{h}{{}^{|j]}{}^{\nu}}\tensor{V}{{}_{\nu}}\Big{]}\Sigma{{}_{ij}}$
$\ \ \ +w\Big{[}\tensor{V}{{}_{\mu}}+{\textstyle\frac{2}{3}}\tensor{h}{{}_{k}^{\lambda}}\tensor{\partial}{{}_{[\lambda|}}\tensor{b}{{}^{k}_{|\mu]}}\Big{]}\Big{)}\varphi.$ (16)
In this $\big{\{}\tensor{b}{{}^{i}_{\mu}},\tensor{\tilde{A}}{{}^{ij}_{\mu}},\tensor{V}{{}_{\mu}},\phi\big{\}}$ formulation $\tensor{V}{{}_{\mu}}$ is a ‘Lorentz’ boson, not a ‘Weyl’ boson, because $\tensor{V}{{}_{\mu}}\mapsto\tensor{V}{{}_{\mu}}-\frac{1}{3}\tensor{h}{{}^{\lambda}_{l}}\tensor{b}{{}^{m}_{\mu}}\tensor{\Lambda}{{}_{k}^{l}}\tensor{\partial}{{}_{\lambda}}\tensor{\Lambda}{{}^{k}_{m}}$. Relabelling $\tensor{A}{{}^{ij}_{\mu}}\equiv\tensor{\tilde{A}}{{}^{ij}_{\mu}}-2\tensor{b}{{}^{[i|}_{\mu}}\tensor{h}{{}^{|j]}{}^{\nu}}\tensor{V}{{}_{\nu}}$, where $\tensor{A}{{}^{ij}_{\mu}}\equiv\tensor{A}{{}^{[ij]}_{\mu}}$ carries 24 d.o.f, ensures the PGT transformation of $\tensor{A}{{}^{ij}_{\mu}}$. In this $\big{\{}\tensor{b}{{}^{i}_{\mu}},\tensor{A}{{}^{ij}_{\mu}},\phi\big{\}}$ formulation Eq. 16 becomes
$\tensor{\mathscr{D}}{{}_{\mu}}\varphi$ (+$\equiv\Big{(}\tensor{\partial}{{}_{\mu}}+{\textstyle\frac{1}{2}}\tensor{A}{{}^{ij}_{\mu}}\Sigma{{}_{ij}}+{\textstyle\frac{1}{3}}w\tensor{h}{{}_{k}^{\lambda}}\Big{[}2\tensor{\partial}{{}_{[\lambda|}}\tensor{b}{{}^{k}_{|\mu]}}+3\tensor{b}{{}_{l}{}_{\mu}}\tensor{A}{{}^{kl}_{\lambda}}\Big{]}\Big{)}\varphi$
(+$\equiv\Big{(}\tensor{\partial}{{}_{\mu}}+{\textstyle\frac{1}{2}}\tensor{A}{{}^{ij}_{\mu}}\Sigma{{}_{ij}}+{\textstyle\frac{1}{3}}w\tensor{T}{{}_{\mu}}\Big{)}\varphi,$ (17)
where we used Eq. 10a in the last equality. New fields should not be introduced where they are not needed: in hindsight it was obvious from Eq. 12 that PGT contained a vector $\tensor{T}{{}_{\mu}}/3$ which already performed the function of $\tensor{B}{{}_{\mu}}$. This fact was observed by Obukhov in [187], and the covariant derivative in Eq. 17 was first written down in [156] (see also [188]). One could interpret Eq. 17 as a hint that PGT descends from a scale-invariant theory. Explicit scales emerge in the $\big{\{}\tensor{\hat{b}}{{}^{i}_{\mu}},\tensor{\hat{A}}{{}^{ij}_{\mu}}\big{\}}$ formulation, with scale-invariant variables $\tensor{b}{{}^{i}_{\mu}}\mapsto\phi_{0}\tensor{\hat{b}}{{}^{i}_{\mu}}/\phi$ and $\tensor{A}{{}^{ij}_{\mu}}\mapsto\tensor{\hat{A}}{{}^{ij}_{\mu}}$. By dropping the hats on variables, we find that the PGT Lagrangian couplings parameterise exactly the same physics as the following combinations of eWGT couplings (with $\alpha_{1},\ldots,\alpha_{6}$ identical)
$\begin{gathered}{m_{\mathrm{P}}}{}^{2}\equiv\zeta\phi_{0}^{2},\quad{m_{\mathrm{P}}}{}^{2}\Lambda\equiv\lambda\phi_{0}^{4},\quad\tensor{\beta}{{}_{1}}\equiv\tensor{\gamma}{{}_{1}}\phi_{0}^{2},\\ \tensor{\beta}{{}_{2}}\equiv\tensor{\gamma}{{}_{2}}\phi_{0}^{2},\quad\tensor{\beta}{{}_{3}}\equiv\big{(}\nu-6(2\tensor{\gamma}{{}_{1}}+\tensor{\gamma}{{}_{2}})\big{)}\phi_{0}^{2}/18.\end{gathered}$ (18)
Although the $\tensor{\gamma}{{}_{3}}$ coupling of WGT in Weyl gauge theory (WGT) is missing from the eWGT action, its loss is exactly compensated for by $\nu$ so that $\tensor{\beta}{{}_{3}}$ emerges in Eq. 18 as an independent coupling in PGT. The couplings $\chi$ and $\xi$ do not appear in Eq. 18 — these parameterise terms involving $\tensor{H}{{}_{\mu\nu}}$, which is the Faraday tensor for $\tensor{T}{{}_{\mu}}$ in the scale-invariant variables of PGT. Thus, eWGT motivates a correction to Yang–Mills-type PGT:
$S_{\text{eWGT}}\equiv S_{\text{PGT}}+\frac{1}{3}\int\mathrm{d}^{4}x\sqrt{-g}\Big{[}$ $\chi\tensor{R}{{}_{[\mu\nu]}}\tensor{\partial}{{}^{[\mu}}\tensor{T}{{}^{\nu]}}$
$-\frac{\xi}{12}\tensor{\partial}{{}_{[\mu}}\tensor{T}{{}_{\nu]}}\tensor{\partial}{{}^{[\mu}}\tensor{T}{{}^{\nu]}}\Big{]}.$ (19)
[…]
Implications for the PGT spectrum
Since we have made a substantial correction to the PGT Lagrangian, it is important to update the well-known PGT particle spectrum in [60]. To do this, we take $\lambda=\Lambda=0$ and focus on the linearisation near Minkowski spacetime. As shown in Appendix A and Fig. 4, the propagator of $S_{\text{eWGT}}$ in Equivalence of PGT and eWGT contains a quartic pole in the parity-odd spin-one sector
$2\Big{[}\big{(}2\tensor{\alpha}{{}_{2}}+4\tensor{\alpha}{{}_{4}}+\tensor{\alpha}{{}_{5}}\big{)}\xi-\chi^{2}\Big{]}k^{4}+3\Big{[}96\tensor{\alpha}{{}_{2}}\big{(}2\tensor{\beta}{{}_{1}}+\tensor{\beta}{{}_{2}}+\tensor{\beta}{{}_{3}}\big{)}$
$+8\big{[}24\tensor{\alpha}{{}_{4}}\big{(}2\tensor{\beta}{{}_{1}}+\tensor{\beta}{{}_{2}}+\tensor{\beta}{{}_{3}}\big{)}+6\tensor{\alpha}{{}_{5}}\big{(}2\tensor{\beta}{{}_{1}}+\tensor{\beta}{{}_{2}}+\tensor{\beta}{{}_{3}}\big{)}$
$+\chi\big{(}4\tensor{\beta}{{}_{1}}+2\tensor{\beta}{{}_{2}}-{m_{\mathrm{P}}}{}^{2}\big{)}\big{]}+\big{(}4\tensor{\beta}{{}_{1}}+2\tensor{\beta}{{}_{2}}-{m_{\mathrm{P}}}{}^{2}\big{)}\xi\Big{]}k^{2}$
$+72\big{(}4\tensor{\beta}{{}_{1}}+2\tensor{\beta}{{}_{2}}-{m_{\mathrm{P}}}{}^{2}\big{)}\big{(}2\tensor{\beta}{{}_{1}}+\tensor{\beta}{{}_{2}}+3\tensor{\beta}{{}_{3}}+{m_{\mathrm{P}}}{}^{2}\big{)}.$ (20)
The pole takes the familiar $k^{2}-M^{2}$ form when the $k^{4}$ coefficient is removed from Implications for the PGT spectrum. For example, the branch $2\tensor{\alpha}{{}_{2}}+4\tensor{\alpha}{{}_{4}}+\tensor{\alpha}{{}_{5}}=\chi=0$ contains the Einstein–Proca theory (EPT) $S_{\text{EPT}}\equiv\int\mathrm{d}^{4}x\sqrt{-g}\big{[}-\frac{1}{2}{m_{\mathrm{P}}}{}^{2}\tensor{\mathring{R}}{}-\frac{1}{4}\tensor{\partial}{{}_{[\mu}}\tensor{T}{{}_{\nu]}}\tensor{\partial}{{}^{[\mu}}\tensor{T}{{}^{\nu]}}+\frac{1}{2}M^{2}\tensor{T}{{}_{\mu}}\tensor{T}{{}^{\mu}}+\text{matter}\big{]}$ in $\big{\{}\tensor{g}{{}_{\mu\nu}},\tensor{T}{{}^{\alpha}_{\mu}{}_{\nu}}\big{\}}$ variables. EPT is evidently not strongly coupled, so this is the much-sought-after vector torsion. The more general branch $\tensor{\alpha}{{}_{5}}=\chi^{2}/\xi-2\left(\tensor{\alpha}{{}_{2}}+2\tensor{\alpha}{{}_{4}}\right)$ leads to the new mass spectrum in Eqs. 24a to 24f, as shown in Fig. 5. It is also interesting to instead remove the constant term in Implications for the PGT spectrum. This produces a $k^{2}\big{(}k^{2}-M^{2}\big{)}$ pole, which affects the massless spectrum. As shown in Fig. 6, we find an example case $\tensor{\alpha}{{}_{2}}=\tensor{\alpha}{{}_{3}}-4\tensor{\alpha}{{}_{4}}+\tensor{\alpha}{{}_{6}}=4\tensor{\alpha}{{}_{4}}+\tensor{\alpha}{{}_{5}}=4\tensor{\beta}{{}_{1}}+2\tensor{\beta}{{}_{2}}-{m_{\mathrm{P}}}{}^{2}=0$ which propagates (without ghosts or tachyons) the Einstein graviton, one heavy scalar, one heavy pseudoscalar, and a one extra massless scalar. To understand where this scalar comes from, we strip away all the uninvolved operators to show in Fig. 7 that the model
$S=\int\mathrm{d}^{4}x\sqrt{-g}\big{[}\chi\tensor{R}{{}_{[\mu\nu]}}\tensor{\partial}{{}^{[\mu}}\tensor{T}{{}^{\nu]}}+\tensor{\beta}{{}_{3}}\tensor{T}{{}_{\mu}}\tensor{T}{{}^{\mu}}\big{]},$ (21)
propagates, for $\chi\neq 0$, one massless scalar with the no-ghost condition $\tensor{\beta}{{}_{3}}>0$. For $\chi=0$, the spectrum is empty. In fact Eq. 21 can be reduced even further by removing $\tensor{b}{{}^{i}_{\mu}}$, and keeping only the axial-free part of $\tensor{A}{{}^{ij}_{\mu}}$. This part has the index symmetries of a Curtright field $\tensor{C}{{}_{\mu\nu\sigma}}$, i.e. $\tensor{C}{{}_{\mu\nu\sigma}}\equiv\tensor{C}{{}_{[\mu\nu]\sigma}}$ with $\tensor{C}{{}_{[\mu\nu\sigma]}}\equiv 0$ [190], and Eq. 21 reduces (linearly) to
$S=\int\mathrm{d}^{4}x\Big{[}\tensor{\beta}{{}_{3}}\tensor{C}{{}^{\mu}}\tensor{C}{{}_{\mu}}+\frac{\chi}{12}\tensor{C}{{}_{\mu\nu}}\Big{(}\tensor{C}{{}^{\mu\nu}}-\frac{4}{3}\tensor{\partial}{{}_{\sigma}}\tensor{C}{{}^{\mu(\nu\sigma)}}\Big{)}\Big{]},$ (22)
with trace $\tensor{C}{{}_{\mu}}\equiv\tensor{C}{{}^{\nu}_{\mu\nu}}$ and field strength $\tensor{C}{{}_{\mu\nu}}\equiv 2\tensor{\partial}{{}_{[\mu}}\tensor{C}{{}_{\nu]}}$. As shown in Fig. 8Eqs. 22 and 21 have the same spectra. Finally, in the branch $\xi=\chi=0$ we recover the known spectrum of $S_{\text{PGT}}$ in No-go theorem for vector torsion, as first presented in [60].
Closing remarks
Poincaré gauge theory (PGT) strongly motivates spacetime torsion [2, 3, 4], but the consensus since the turn of the millennium has been that PGT prohibits vector torsion from propagating [140, 141, 61, 62, 70, 71, 74, 76, 77, 78, 82, 86, 89, 91, 97, 101]. Separately, the full conformal group is a more appealing gauge group than the Poincaré group. With this context, our letter achieved three objectives:
  1. 1.
    Starting just with the inflaton mass, scale-invariant embedding leads to the slow-roll plateau in Eqs. 8b and 1. This is an intriguing, stand-alone result.
  2. 2.
    We showed that PGT has a unique, locally scale-invariant embedding as extended Weyl gauge theory (eWGT) [156, 157, 158, 159, 160]. The embedding is unique if one is to achieve scale invariance via the introduction of a minimum number of new gauge fields, beyond those already present in PGT. To reach eWGT, one need only introduce a compensator field. The compensator is purely gauge, so that the embedding theory is completely indistinguishable from PGT after gauge-fixing. In other words, PGT already descends from a conformal theory, without needing further work. By contrast, to reach the traditional Weyl gauge theory (WGT) embedding one must introduce both a compensator and a Weyl gauge boson. WGT actions may be written down which propagate this Weyl boson even after gauge fixing resulting in phenomena that cannot be described by PGT alone.
  3. 3.
    The eWGT embedding of PGT means that the natural Yang–Mills-type PGT action is missing the two terms in Equivalence of PGT and eWGT. The new terms provide the first compelling argument for propagating vector torsion. The updated mass spectrum of PGT is shown in Eqs. 24a to 24f. The new terms reveal that the Curtright field in Eq. 22 is embedded in the PGT dynamics.
Further work is needed to develop our inflaton model in the full eWGT-PGT framework. One key question is whether the model is susceptible to the Weyl anomaly [191]. Also, the implications of the new PGT terms for strong coupling should be investigated [129]. Finally, the Yang–Mills-type actions in No-go theorem for vector torsion and Weyl gauge theory (WGT) are restricted to parity-even terms for simplicity: the parity-odd extensions should be considered. The mass spectrum of parity-violating PGT was found in [64], and confirmed in [65], however the massless spectra and unitarity conditions of the various critical cases of the theory have not been thoroughly explored, nor have more than a handful of such cases been identified to date. [^16] This is due to the fact that, in the parity-violating case, the formulae for the squares of the masses are irrational functions of the couplings: the residues of such poles are comparatively more cumbersome to calculate. As a result, the capacity for parity-violating PGT to propagate a vector from Yang–Mills-type terms has not been fully explored linearly, and the question of strong coupling in such models has not, to our knowledge, been addressed at all.
Acknowledgements.
We are grateful for several useful discussions with Georgios Karananas, Carlo Marzo, Prabhoda Chandra Sarjapur and Sebastian Zell, and for the feedback of the anonymous referee. WB is grateful for the kind hospitality of Leiden University and the Lorentz Institute, and the support of Girton College, Cambridge. Y-CL acknowledges support from the Ministry of Education of Taiwan and the Cambridge Commonwealth, European & International Trust via a Taiwan Cambridge Scholarship, and from Corpus Christi College, Cambridge. This work used the DiRAC Data Intensive service (CSD3 www.csd3.cam.ac.uk) at the University of Cambridge, managed by the University of Cambridge University Information Services on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The DiRAC component of CSD3 at Cambridge was funded by BEIS, UKRI and STFC capital funding and STFC operations grants. DiRAC is part of the UKRI Digital Research Infrastructure. This work also used the Newton server, to which access was provided by Will Handley, and funded through an ERC grant.
Appendix A Detailed tree-level spectra
Figure 2: Kinematic structure of the tetrad perturbation $\tensor{f}{{}_{\mu\nu}}$, as used in No-go theorem for vector torsion and Equivalence of PGT and eWGT. These definitions are used in Figs. 4 to 11. See [192] for further notational details. This is a vector graphic: all details are visible under magnification.
Figure 3: Kinematic structure of the spin connection perturbation $\tensor{A}{{}_{\mu\nu\sigma}}$, as used in No-go theorem for vector torsion and Equivalence of PGT and eWGT. Note that the $2^{-}$ state has a hidden multi-term cyclic symmetry on all its indices, which is not accommodated by the C language implementation of the Butler–Portugal algorithm [193]. These definitions are used in Figs. 4 to 11. See [192] for further notational details. This is a vector graphic: all details are visible under magnification.
The particle spectra are obtained using the Particle Spectrum for Any Tensor Lagrangian (PSALTer) software [194, 106, 177, 192]. Recall that the kinematic variables of PGT were $\big{\{}\tensor{b}{{}^{i}_{\mu}},\tensor{A}{{}^{ij}_{\mu}}\big{\}}$, and that the inverse $\tensor{h}{{}_{i}^{\mu}}$ is determined by $\tensor{b}{{}^{i}_{\mu}}$. This means that we are free to use $\big{\{}\tensor{h}{{}_{i}^{\mu}},\tensor{A}{{}^{ij}_{\mu}}\big{\}}$ instead. Working in the weak-field regime, we take $\tensor{A}{{}^{ij}_{\mu}}$ to be inherently perturbative. We define the exact perturbation $\tensor{h}{{}_{i}^{\mu}}\equiv\tensor{*}{\delta}{{}_{i}^{\mu}}+\tensor{f}{{}_{i}^{\mu}}$, where we make the ‘Kronecker’ choice of Minkowski vacuum (see alternative vacua [194, 195, 196]). There are a priori 16 d.o.f in $\tensor{f}{{}_{i}^{\mu}}$, and 24 d.o.f in $\tensor{A}{{}^{ij}_{\mu}}$. Conjugate to $\tensor{f}{{}_{i}^{\mu}}$ and $\tensor{A}{{}^{ij}_{\mu}}$ are the translational source (i.e. the asymmetric stress-energy tensor) $\tensor{\tau}{{}^{i}_{\mu}}$, and the matter spin current $\tensor{\sigma}{{}_{ij}^{\mu}}$ [18, 46, 17]. To lowest order in the field perturbations, the Greek and Roman indices are interchangeable. For this reason the final set of fields used in the linearised analysis is $\big{\{}\tensor{f}{{}_{\mu\nu}},\tensor{A}{{}_{\mu\nu\sigma}}\big{\}}$, with conjugate sources $\big{\{}\tensor{\tau}{{}^{\mu\nu}},\tensor{\sigma}{{}^{\mu\nu\sigma}}\big{\}}$. The various $\mathrm{SO}(3)$ irreducible parts of these quantities are presented in Figs. 2 and 3, and they have spin-parity ($J^{P}$) labels to identify them. Duplicate $J^{P}$ states can arise, and these are distinguished by extra labels #$\#1$#$\#2$, etc.
The actual analysis of the theory in Equivalence of PGT and eWGT was performed across 64 AMD® Ryzen Threadripper CPUs. For the completely unrestricted case in Equivalence of PGT and eWGT the mass expressions associated with the roots of the quartic pole are expected to be very cumbersome, so we abort the computation before they are computed: the structure of the wave operator and saturated propagator, without the mass spectrum, is shown in Fig. 4. Before the evaluation of the ‘general’ branch for which the $k^{4}$ coefficient in Implications for the PGT spectrum is removed, the condition $\tensor{\alpha}{{}_{5}}=\chi^{2}/\xi-2\left(\tensor{\alpha}{{}_{2}}+2\tensor{\alpha}{{}_{4}}\right)$ must be imposed on the linearised action. To avoid Lagrangian couplings appearing on the denominator of the linearised action [^17] As a mild technical limitation, coupling coefficients in the denominator can cause PSALTer to become slow. we impose the constraint by introducing a new coupling $\theta$ and setting in Equivalence of PGT and eWGT
$\tensor{\alpha}{{}_{5}}\mapsto\theta^{2}\xi-2\left(\tensor{\alpha}{{}_{2}}+2\tensor{\alpha}{{}_{4}}\right),\quad\chi\mapsto\theta\xi.$ (23)
The results are shown in Fig. 5, we see that apart from the two massless polarisations of the Einstein graviton, up to six torsion particles can propagate with the following square masses, where we eliminate $\theta$ in terms of $\chi$ and $\xi$ using Eq. 23
$\left(\tensor{m}{{}_{0^{+}}}\right)^{2}$ $\equiv\frac{{m_{\mathrm{P}}}{}^{2}\left(2\tensor{\beta}{{}_{1}}+\tensor{\beta}{{}_{2}}+3\tensor{\beta}{{}_{3}}+{m_{\mathrm{P}}}{}^{2}\right)}{2\left(2\tensor{\beta}{{}_{1}}+\tensor{\beta}{{}_{2}}+3\tensor{\beta}{{}_{3}}\right)\left(6\tensor{\alpha}{{}_{1}}+2\tensor{\alpha}{{}_{3}}-2\tensor{\alpha}{{}_{4}}+2\tensor{\alpha}{{}_{6}}+\chi^{2}/\xi\right)},$ (24a)
$\left(\tensor{m}{{}_{0^{-}}}\right)^{2}$ $\equiv-\frac{8\tensor{\beta}{{}_{1}}-8\tensor{\beta}{{}_{2}}+{m_{\mathrm{P}}}{}^{2}}{4\tensor{\alpha}{{}_{2}}+12\tensor{\alpha}{{}_{4}}-2\chi^{2}/\xi},$ (24b)
$\left(\tensor{m}{{}_{1^{+}}}\right)^{2}$ $\equiv\frac{-32\tensor{\beta}{{}_{1}}^{2}+16\tensor{\beta}{{}_{2}}^{2}-10\tensor{\beta}{{}_{2}}{m_{\mathrm{P}}}{}^{2}+{m_{\mathrm{P}}}{}^{4}+4\tensor{\beta}{{}_{1}}\left(4\tensor{\beta}{{}_{2}}+{m_{\mathrm{P}}}{}^{2}\right)}{4\left(\tensor{\alpha}{{}_{2}}-\tensor{\alpha}{{}_{3}}+4\tensor{\alpha}{{}_{4}}-4\tensor{\alpha}{{}_{6}}\right)\left(2\tensor{\beta}{{}_{1}}-\tensor{\beta}{{}_{2}}\right)},$ (24c)
$\left(\tensor{m}{{}_{1^{-}}}\right)^{2}$ $\equiv-\frac{24\left(4\tensor{\beta}{{}_{1}}+2\tensor{\beta}{{}_{2}}-{m_{\mathrm{P}}}{}^{2}\right)\left(2\tensor{\beta}{{}_{1}}+\tensor{\beta}{{}_{2}}+3\tensor{\beta}{{}_{3}}+{m_{\mathrm{P}}}{}^{2}\right)}{\xi\left(48\tensor{\beta}{{}_{3}}\chi^{3}/\xi^{3}+4\tensor{\beta}{{}_{1}}\left(1+8\chi/\xi+24\chi^{2}/\xi^{2}\right)+2\tensor{\beta}{{}_{2}}\left(1+8\chi/\xi+24\chi^{2}/\xi^{2}\right)-{m_{\mathrm{P}}}{}^{2}-8{m_{\mathrm{P}}}{}^{2}\chi/\xi\right)},$ (24d)
$\left(\tensor{m}{{}_{2^{+}}}\right)^{2}$ $\equiv\frac{\left(4\tensor{\beta}{{}_{1}}+2\tensor{\beta}{{}_{2}}-{m_{\mathrm{P}}}{}^{2}\right){m_{\mathrm{P}}}{}^{2}}{4\left(2\tensor{\beta}{{}_{1}}+\tensor{\beta}{{}_{2}}\right)\left(3\tensor{\alpha}{{}_{2}}-\tensor{\alpha}{{}_{3}}+4\tensor{\alpha}{{}_{4}}-4\tensor{\alpha}{{}_{6}}-2\chi^{2}/\xi\right)},$ (24e)
$\left(\tensor{m}{{}_{2^{-}}}\right)^{2}$ $\equiv\frac{4\tensor{\beta}{{}_{1}}+2\tensor{\beta}{{}_{2}}-{m_{\mathrm{P}}}{}^{2}}{4\tensor{\alpha}{{}_{2}}-2\chi^{2}/\xi}.$ (24f)
By adding up the $1+1+3+3+5+5$ spin multiplicities of Eqs. 24a to 24f we recover the $16+24$ kinematic d.o.f in $\big{\{}\tensor{b}{{}^{i}_{\mu}},\tensor{A}{{}^{ij}_{\mu}}\big{\}}$, less the $2\times(4+6)$ gauge d.o.f associated with Poincaré gauge symmetry — the only surviving symmetry of the embedded theory. The branch $2\tensor{\alpha}{{}_{2}}+4\tensor{\alpha}{{}_{4}}+\tensor{\alpha}{{}_{5}}=\chi=0$ is likewise found by setting $\tensor{\alpha}{{}_{5}}\mapsto-2\left(\tensor{\alpha}{{}_{2}}+2\tensor{\alpha}{{}_{4}}\right)$ and $\chi\mapsto 0$ in Equivalence of PGT and eWGT. The analysis is presented in Fig. 9, and shows that the mass spectrum is fully consistent with imposing $\chi\mapsto 0$ in Eqs. 24a to 24f. Finally, the case $\xi=\chi=0$ in Equivalence of PGT and eWGT without any further restrictions leads in Fig. 10 to the general case of PGT in No-go theorem for vector torsion, which was found previously in [60], namely
$\left(\tensor{m}{{}_{0^{+}}}\right)^{2}$ $\equiv\frac{{m_{\mathrm{P}}}{}^{2}\left(2\tensor{\beta}{{}_{1}}+\tensor{\beta}{{}_{2}}+3\tensor{\beta}{{}_{3}}+{m_{\mathrm{P}}}{}^{2}\right)}{2\left(6\tensor{\alpha}{{}_{1}}+2\tensor{\alpha}{{}_{2}}+2\tensor{\alpha}{{}_{3}}+2\tensor{\alpha}{{}_{4}}+\tensor{\alpha}{{}_{5}}+2\tensor{\alpha}{{}_{6}}\right)\left(2\tensor{\beta}{{}_{1}}+\tensor{\beta}{{}_{2}}+3\tensor{\beta}{{}_{3}}\right)},$ (25a)
$\left(\tensor{m}{{}_{0^{-}}}\right)^{2}$ $\equiv-\frac{8\tensor{\beta}{{}_{1}}-8\tensor{\beta}{{}_{2}}+{m_{\mathrm{P}}}{}^{2}}{4\tensor{\alpha}{{}_{4}}-2\tensor{\alpha}{{}_{5}}},$ (25b)
$\left(\tensor{m}{{}_{1^{-}}}\right)^{2}$ $\equiv-\frac{\left(4\tensor{\beta}{{}_{1}}+2\tensor{\beta}{{}_{2}}-{m_{\mathrm{P}}}{}^{2}\right)\left(2\tensor{\beta}{{}_{1}}+\tensor{\beta}{{}_{2}}+3\tensor{\beta}{{}_{3}}+{m_{\mathrm{P}}}{}^{2}\right)}{2\left(2\tensor{\alpha}{{}_{2}}+4\tensor{\alpha}{{}_{4}}+\tensor{\alpha}{{}_{5}}\right)\left(2\tensor{\beta}{{}_{1}}+\tensor{\beta}{{}_{2}}+\tensor{\beta}{{}_{3}}\right)},$ (25c)
$\left(\tensor{m}{{}_{2^{+}}}\right)^{2}$ $\equiv\frac{{m_{\mathrm{P}}}{}^{2}\left(-4\tensor{\beta}{{}_{1}}-2\tensor{\beta}{{}_{2}}+{m_{\mathrm{P}}}{}^{2}\right)}{4\left(\tensor{\alpha}{{}_{2}}+\tensor{\alpha}{{}_{3}}+4\tensor{\alpha}{{}_{4}}+2\tensor{\alpha}{{}_{5}}+4\tensor{\alpha}{{}_{6}}\right)\left(2\tensor{\beta}{{}_{1}}+\tensor{\beta}{{}_{2}}\right)},$ (25d)
$\left(\tensor{m}{{}_{2^{-}}}\right)^{2}$ $\equiv\frac{-4\tensor{\beta}{{}_{1}}-2\tensor{\beta}{{}_{2}}+{m_{\mathrm{P}}}{}^{2}}{2\left(4\tensor{\alpha}{{}_{4}}+\tensor{\alpha}{{}_{5}}\right)}.$ (25e)
Whilst we confirmed already that the $\chi\to 0$ limit of Eqs. 24a to 24f was continuous, it is evident from Eq. 24d that once $\chi\to 0$ has been taken the limit $\xi\to 0$ eliminates the $1^{-}$ vector in which we have been so interested. In fact, this is not surprising, because taking $\chi\to 0$ at finite $\xi$ corresponds to setting $2\tensor{\alpha}{{}_{2}}+4\tensor{\alpha}{{}_{4}}+\tensor{\alpha}{{}_{5}}=0$. From Eq. 25c we see how this latter condition would equivalently kill the $1^{-}$ mode in No-go theorem for vector torsion (we confirm this in Fig. 11), and the No-go theorem for vector torsion theory-space is reached by $\xi\to 0$.
Figure 4: Partial particle spectrograph of the completely general theory in Equivalence of PGT and eWGT. From the $1^{-}$ sector of the saturated propagator we may read off the quartic pole in Implications for the PGT spectrum. All quantities are defined in Figs. 2 and 3. See [192] for further notational details. This is a vector graphic: all details are visible under magnification.
Figure 5: Particle spectrograph of the branch $\tensor{\alpha}{{}_{5}}=\chi^{2}/\xi-2\left(\tensor{\alpha}{{}_{2}}+2\tensor{\alpha}{{}_{4}}\right)$ of Equivalence of PGT and eWGT, which gives rise to the mass spectra in Eqs. 24a to 24f. All quantities are defined in Figs. 2, 3 and 23. See [192] for further notational details. This is a vector graphic: all details are visible under magnification.
Figure 6: Particle spectrograph of the branch $\tensor{\alpha}{{}_{2}}=\tensor{\alpha}{{}_{3}}-4\tensor{\alpha}{{}_{4}}+\tensor{\alpha}{{}_{6}}=4\tensor{\alpha}{{}_{4}}+\tensor{\alpha}{{}_{5}}=4\tensor{\beta}{{}_{1}}+2\tensor{\beta}{{}_{2}}-{m_{\mathrm{P}}}{}^{2}=0$ of Equivalence of PGT and eWGT. All quantities are defined in Figs. 2 and 3. See [192] for further notational details. This is a vector graphic: all details are visible under magnification.
Figure 7: Particle spectrograph of Eq. 21, i.e. the minimal PGT realisation of the novel massless scalar particle which appears in Fig. 6. All quantities are defined in Figs. 2 and 3. See [192] for further notational details. This is a vector graphic: all details are visible under magnification.
Figure 8: Particle spectrograph of Eq. 22, i.e. the minimal model of the novel scalar particle which appears in Figs. 6 and 7. We use the rotational gauge field perturbation $\tensor{A}{{}_{\mu\nu\sigma}}\equiv\tensor{A}{{}_{[\mu\nu]\sigma}}$ as a proxy for the Curtright field $\tensor{C}{{}_{\mu\nu\sigma}}\equiv\tensor{C}{{}_{[\mu\nu]\sigma}}$ with $\tensor{C}{{}_{[\mu\nu\sigma]}}\equiv 0$, because the latter multi-term symmetry is a difficult attribute to declare in the PSALTer framework. It can be confirmed (by applying transformation rules to the Lagrangian density in this figure, and repeating the analysis) that the axial part $\tensor{A}{{}_{[\mu\nu\sigma]}}$ does not contribute to the spectrum. All quantities are defined in Figs. 2 and 3. See [192] for further notational details. This is a vector graphic: all details are visible under magnification.
Figure 9: Particle spectrograph of the branch $\tensor{\alpha}{{}_{5}}=-2\left(\tensor{\alpha}{{}_{2}}+2\tensor{\alpha}{{}_{4}}\right)$ with $\chi=0$ of Equivalence of PGT and eWGT, which gives rise to the $\chi\mapsto 0$ limit of the mass spectra in Eqs. 24a to 24f (see Fig. 5). All quantities are defined in Figs. 2, 3 and 23. See [192] for further notational details. This is a vector graphic: all details are visible under magnification.
Figure 10: Particle spectrograph of the branch $\xi=\chi=0$ of Equivalence of PGT and eWGT, which gives rise to the mass spectra in Eqs. 25a to 25e, as first presented in [60]. All quantities are defined in Figs. 2 and 3. See [192] for further notational details. This is a vector graphic: all details are visible under magnification.
Figure 11: Particle spectrograph of the branch $\tensor{\alpha}{{}_{5}}=-2\left(\tensor{\alpha}{{}_{2}}+2\tensor{\alpha}{{}_{4}}\right)$ of No-go theorem for vector torsion, confirming that the $1^{-}$ mode is removed. All quantities are defined in Figs. 2 and 3. See [192] for further notational details. This is a vector graphic: all details are visible under magnification.